Opened 8 years ago

Closed 7 years ago

Last modified 20 months ago

#392 closed flight processing (fixed)

EUFAR11/04, flight day 097/2011, HYMEDECOS (Gradients)

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2011/flight_data/spain_portugal/EUFAR11_04-2011_097_HYMEDECOS-gradients/

Data arrived from ARSF via network transfer on 19/04/11.

Scientific objective: Applied studies relating to soil erosion, biodiversity - vegetation and steppe bird - dynamics, estimates of forage crop quality and carbon storage potential will be carried out. Two spatial resolutions.

Study location: South Portugal

Priority: a4h

PI: Patrick Hostert
Co-PI: Patrick Osborne

Other contacts:
Pedro Leitao

Sensors:

  • Eagle (03/08/11)
  • Hawk (03/08/11)
  • Leica LIDAR Replacement (30/08/11)
  • RCD (03/08/11)

Change History (37)

comment:1 Changed 8 years ago by emca

The LiDAR lines at higher altitude (lines 7-12) have very low intensity values - initally thought to be caused by the intemittent AGC issue that affected the calibration data (#389). Queried with Leica and they think it is a result of the very high altitude of flight.

comment:2 Changed 8 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Taking HS.

comment:3 Changed 8 years ago by knpa

The only basestation data for this flight is the ARSF temporary basestation over 100km away from main site. I've also found a EUREF permanent basestation (LAGO) but this is only 30s interval. Temporarily on hold whilst Ops get back to us to let us know if there is any better data for us to use.

comment:4 Changed 8 years ago by anch

  • Description modified (diff)

comment:5 Changed 8 years ago by anch

Processing RAW camera to tif.

comment:6 Changed 8 years ago by anch

Produced thumbnails from TIFFs

comment:7 Changed 8 years ago by knpa

  • Milestone set to 2011 data processing completion

Basestation data received today (05/07/11) from Pedro Leitao. Should be able to unblock this project.

comment:8 Changed 8 years ago by knpa

Was on hold for a few more days whilst we waited for precise basestation coords from Pedro.

anch is now doing the navigation processing.

comment:9 Changed 8 years ago by anch

Navigation has been processed and verified against ASTER to be correct.

comment:10 Changed 8 years ago by anch

Camera ready for delivery check at:

/data/okhotsk1/scratch/arsf/arsf_data/2011/flight_data/spain_portugal/EUFAR11_04-2011_097_HYMEDECOS-gradients/processing/delivery/EUFAR11_04-097-camera-20110718

comment:11 Changed 8 years ago by knpa

  • Description modified (diff)

comment:12 Changed 8 years ago by knpa

  • Description modified (diff)

comment:13 Changed 8 years ago by knpa

Created las files.

Whilst finding height differences/checking for roll, found that the high flightlines are on average ~12cm above the low flightlines. The hight flightlines are particularly high, ~14,000ft above the terrain.

comment:14 Changed 8 years ago by knpa

A possible cause for the above has been identified: the pulse rate changed between the high and low lines, which put it under the TPR. However the IBRC is only single table and the TPR offset -0.001 so there is a question mark over whether the increased pulse width was dealt with properly. Emca is going to query this with leica.

comment:15 Changed 8 years ago by knpa

  • Description modified (diff)

comment:16 Changed 8 years ago by knpa

Found HS scts:

e h
1 -0.06 -0.07
2 -0.04 -0.05
3 -0.03 -0.06
4 -0.02 -0.03
5 -0.01 -0.06
6 -0.05 -0.05
7 -0.01 -0.06
8 -0.05 -0.06
9 -0.02 0.24
10 -0.02 -0.06
11 -0.05 -0.07
12 -0.05 -0.06

comment:17 Changed 8 years ago by knpa

The above were found with aplcorr so needles to say they do not have the -0.055 adjustment.

comment:18 Changed 8 years ago by knpa

Hyperspectral ready for delivery check (EUFAR11_04-097-hyperspectral-20110722)

comment:19 Changed 8 years ago by knpa

Completed RCD Delivery check

No problems found.

This project has a co-pi so may need to be delivered to both of them.

comment:20 Changed 8 years ago by anch

Finished delivery check.

Couple of things:
A few dropped frames at the start of hawk line 8.
Readme is missing some info (e.g. Project code) on first page.

Other than that it's all food and ready to go.

Please note that there is also a camera delivery ready for this project, and it may need to go to a CO PI

comment:21 Changed 8 years ago by knpa

Dropped frames are not severe enough to be mentioned in Readme.

Can't find the missing information in the Readme on page 1 that anch was referring to.

Preparing delivery for postage (along with RCD). With current pulse rate height discrepancy investigation, lidar may take a little longer.

comment:22 Changed 8 years ago by knpa

  • Description modified (diff)

Hyperspectral and RCD delivered today (03/08/11)

Two packages sent, one to Patrick Osborne in Southampton and one to Patrick Hostert in Berlin.

comment:23 Changed 8 years ago by knpa

LIDAR: Leica have sent us a dual IBRC table (SN084_20110509_DUAL_IBRC.csv) which has improved the discrepancy. To remove the remaining difference they suggested adding a non-zero TPR range offset for use with this table. A value of -0.04 appears to give the optimal result.

Unblocking sensor and continuing processing.

comment:24 Changed 8 years ago by knpa

Relatively large discrepancies between the high alt flightlines which is probably to be expected. No consistent roll issue identified in project but one line (13:08, 005_45a) has a massive roll (0.5m-0.8m) when compared to an another line flew later in the same place and same direction (14:43, 005_45b). By comparing it to the other lines that cross it (esp low flightlines which are more accurate), have decided that it is the 13:08 line which is askew so will attempt to correct that to match 14:43.

comment:25 Changed 8 years ago by knpa

Found an approximate fix for 13:08 using a roll value of -0.0032

comment:26 Changed 8 years ago by knpa

Finished QC stage. Preparing delivery.

comment:27 Changed 8 years ago by knpa

LiDAR ready for delivery check at:

~knpa/2011_arsf_data/spain_portugal/EUFAR11_04-2011_097_HYMEDECOS-gradients/processing/delivery/EUFAR11_04-097-lidar-20110819

comment:28 Changed 8 years ago by adbe

Starting Delivery Check

comment:29 Changed 8 years ago by adbe

Removed .xml files from dem directory

Data ready to go

comment:30 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

LiDAR dispatched today (30/08/11) to Patrick Hostert (Berlin) and Patrick Osborne (Southampton) (two packages).

comment:31 Changed 8 years ago by knpa

Beginning archiving.

comment:32 Changed 8 years ago by knpa

Tarball created and made available to NEODC (22/12/11)

comment:33 Changed 8 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:34 Changed 8 years ago by knpa

Archived by NEODC, moved to archive disk.

comment:35 Changed 7 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

comment:36 Changed 20 months ago by asm

Archiving

Mapped files were not present in the hyperspectral delivery. Created the mapped files and a new delivery in order to create as well JSON files for the EUFAR flight finder tool. New delivery and JSON files are being uploaded to CEDA (started on 15/01/2018).

comment:37 Changed 20 months ago by asm

Archiving

Upload finished, CEDA notified on 18/01/2017.

Note: See TracTickets for help on using tickets.